OpenStudy (anonymous):

if the endpoints of the diameter of a cirlce are (3,1) and (6,5) find the length of the diameter

OpenStudy (anonymous):

the length of two points comes from: [(x1-x2)^2 + (y1-y2)^2]^0.5 now find it!

OpenStudy (anonymous):

i didnt learn how to do this

OpenStudy (anonymous):

the length of the diameter is the distance between (3,1) and (6,5) ok as suggested by shining here x1=3 y1=1 and x2=6 and y2=5 thus d =sqrt((3-6)^2 +(1-5)^2))=sqrt( 9+16)=sqrt(25) = 5 units